OpenStudy (zmudz):

Let \(f(x)\) be a function defined for all positive real numbers satisfying the conditions \(f(x) > 0\) for all \(x > 0\) and \(f(x - y) = \sqrt{f(xy) + 1}\) for all \(x > y > 0\). Determine \(f(2009)\).

OpenStudy (jhannybean):

@AravindG

OpenStudy (jhannybean):

Because 1 is a little bigger than 0?...

OpenStudy (jhannybean):

Confused!!!

OpenStudy (dan815):

ah nvm scratch this stuff i dont think its the right way

OpenStudy (beginnersmind):

dan I think the math is just off. Use used f(2) = 2 :)

OpenStudy (dan815):

oh lets use this we can use that induction step after all

OpenStudy (beginnersmind):

The start is to rewrite \(f(x - y) = \sqrt{f(xy) + 1}\) as \[f(xy) = [f(x-y)]^2 - 1\] In particular using that 2009 = 49-41=8 \[f(2009) = [f(8)]^2 - 1\] So we're not looking for a closed formula. Just a few values like f(1) and f(2) would be enough.

OpenStudy (beginnersmind):

edit: I meant 2009 = 49*41 of course.

OpenStudy (jhannybean):

oohhh haha ok that makes a little bit more sense lol

OpenStudy (dan815):